OpenStudy (anonymous):

Help Please :) calculate dy/dx ; y= (x^2.7/8 + 3/x^2.7)(8x-1)

geerky42 (geerky42):

Use Product Rule.

OpenStudy (anonymous):

how would i start it

geerky42 (geerky42):

Derivative first factor and left second one untouched. Then put sun then derivative second factor and left first one untouched. EX: f(x) = (x^2.7/8 + 3/x^2.7) g(x) = (8x-1) dy/dx = f'(x)g(x) + f(x)g'(x)

geerky42 (geerky42):

Is this clear enough?

OpenStudy (anonymous):

yes, im just having trouble finding the derivative of the first portion of the equation.

OpenStudy (anonymous):

.3375x^5.4-8.1 / x^3.7

geerky42 (geerky42):

Not sure how you got x^5.4... While taking the derivative, you basically should go from x^n to nx^(n-1)

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

what should it be? i got 2.1